OpenStudy (anonymous):

can someone show me how this function would be graphed? y = -x + x and there is a little 2 over the last x? please help

OpenStudy (anonymous):

you mean like this: \(\large y=-x+x^2 \) ???

OpenStudy (anonymous):

or this: \(\large y=-x+\frac{2}{x} \) ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

ok... can you factor the right side for me: \(\large y=-x+x^2 \) ???

OpenStudy (anonymous):

having it factored will tell us where the graph will cross the x-axis (the zeros)

OpenStudy (anonymous):

maybe writing the equation this way is better: \(\large y=x^2-x \) factor the right side....

OpenStudy (anonymous):

i dont understand how to do the graph at all

OpenStudy (anonymous):

i have been haveing a hard time with these i need someone to draw the graph and explain how to do it

OpenStudy (anonymous):

that's ok... you're question was to help graph.. i'm trying to do that but you need to answer my question on factoring first.... all these preliminary stuff will help graph the parabola....

OpenStudy (anonymous):

main thing when graphing is to find the x and y intercepts.... factoring the right side will help locate the x intercepts....

OpenStudy (anonymous):

ok how do i find out what x and y means?

OpenStudy (anonymous):

ok... you can factor \(\large y=x^2-x=x(x-1) \), agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the equation can be written as: \(\large y=x(x-1) \) so it's easy to see that when x=0, y=0 what about when x=1, y= ???

OpenStudy (anonymous):

would it be 1

OpenStudy (anonymous):

try again... \(\large y=\color{red}x(\color{red}x-1) \) when x=1: \(\large y=\color{red}1 (\color{red}1-1) \)

OpenStudy (anonymous):

so it would be 0 then right?

OpenStudy (anonymous):

correct....

OpenStudy (anonymous):

we've just found the x-intercepts... meaning the graph will go through these two points

OpenStudy (anonymous):